OpenStudy (anonymous):

125% as a fraction and a decimal

OpenStudy (freckles):

Replace the % with *1/100

OpenStudy (freckles):

or if you want it in decimal replace the % with *.01

OpenStudy (freckles):

example 115% to fraction we get \[115 \cdot \frac{1}{100}=\frac{115}{100} \text{ note: you need \to reduce} =\frac{115 \div 5}{100 \div 5}=\frac{23}{20}\] 115% to decimal we get \[115 \cdot .01 =1.15\]

OpenStudy (arbyj):

125/100=1.25 125/5=25/5=5 100/5=20/5=4 fraction is 5/4